OpenStudy (anonymous):

Suppose a ranger at the top of a 90-foot observation tower sees a fire. The angle of depression to the fire is 3º. What is the horizontal distance to the fire? Round to the nearest foot.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Draw figure, see that tan(3) = 90/s = 90/distance 0.0524= 90/s s = 90/0.0524 = 1718.